# A Noisy Approach to Intrinsically Mixed-State Topological Order

### 20 Mar 2024 v1

@@ I Introduction
Recent years have witnessed remarkable progress in preparing and manipulating such states in programmable quantum simulators [6, 7]. Decoherence is invariably present in these platforms and thus identifying a sharp notion of mixed-state topological order is not merely of fundamental interest, but also of immediate practical import. While any finite temperature is known to destroy topological order (TO) in two spatial dimensions (2D) [8, 9], for local decoherence below a certain threshold, the quantum information encoded in a topological QECC is recoverable. Indeed, the persistence and eventual breakdown of topological order in a pure state |$\ket{\psi}$ subject to a local decoherence channel $\mathcal{E}$ has recently been studied through the lens of the entanglement properties of the “corrupted” density matrix [|$\mathcal{E}[\ket{\psi}\bra{\psi}]$ [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. While naïvely one expects that local errors destroy quantum correlations (and hence TO), the decohered state is not the Gibbs state and can in principle encode structured entanglement. Strikingly, for initial pure symmetry protected topological (SPT) phases, the decohered density matrix can display intrinsically mixed SPT order [20, 21, 22, 23]. While Ref. [24] recently observed a non-trivial topological entanglement negativity in a decohered Toric code (indicating imTO), a systematic framework for addressing intrinsically mixed-state topological order (imTO) is lacking.
@@ III.3 Chiral Semion from Double Semion
As an instance of this mapping, let us take the doubled semion anyon theory as our parent TO. This theory can be realized as a Pauli stabilizer Hamiltonian [29] and its anyons form a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ group under fusion, with elements $\{1,s,\bar{s},s\bar{s}\}$. Here, $s$ is a semion (it has self-statistics $\theta(s)=i$), $\bar{s}$ is an anti-semion ($\theta(\bar{s})=-i$), and $s\bar{s}$ is a boson. Now, we subject a ground state of this system to an error channel that incoherently proliferates the semion $s$ i.e., $\hat{\mathcal{A}}=\{1,s\}$. This corresponds to gauging out $s$, which braids trivially with $\bar{s}$. The resulting anyon theory for the decohered mixed-state is given by $\mathcal{A}=\{1,\bar{s}\}$ i.e., it is the chiral (anti)-semion Abelian anyon theory. This clearly represents an intrinsically mixed-state phase of matter as chiral UMTCs are widely held to be forbidden in the ground states of local gapped Hamiltonians. Since $\bar{s}$ has non-trivial self-statistics, the decohered code space encodes exactly one logical qubit; this example represents the minimal model in which one obtains a chiral anyon theory with an encoded logical qubit.
@@ IV Decoherence as gauging out in general UMTCs
IV.1 $\mathbb{Z}_{2}$ Toric code from Chiral Ising UMTC
@@ IV.3 Non-modular imTOs from Doubled $SU(2)_{k}$ UMTC
Thus, we can obtain a large family of imTOs by exposing the ground states of string-net models to local error channels, where the decohered code space generically retains logical information i.e., it is a decoherence-free subspace. Since we have shown that the resulting imTOs can host transparent anyons, we obtain a partial classification of imTO in terms of braided fusion categories.
@@ VI Discussion
A pressing issue is to characterize our family of imTOs via their entanglement structure. While the entanglement entropy has previously been studied in mixed-state TO [47, 63], a more natural probe of entanglement in mixed-states is provided by the entanglement negativity which, unlike the entanglement entropy, is a good measure of quantum correlations in a mixed state [64, 65, 66, 67, 68]. In pure state TO, the negativity has been shown to receive universal contributions which are sensitive to the modular data of the TO (namely, the total quantum dimension) [69, 70, 71, 72, 73, 74, 9]. This topological entanglement negativity (TEN) has also been shown to be sensitive to the breakdown of TO in thermal states [75, 9], which the entanglement entropy does not accurately reflect. Since we have shown that imTO is generally characterized by non-modular anyon theories, it is an intriguing question as to what universal data the TEN captures in these states. In one specific instance, Ref. [24] distinguished between the $\mathbb{Z}_{2}^{(0)}$ and $\mathbb{Z}_{2}^{(1)}$ imTOs (obtained via decoherence of the $\mathbb{Z}_{2}$ Toric code) by the respective absence and presence of topological contributions to the negativity. Recalling that these two states correspond to quantum condensates of bosonic and fermionic loops, it is tempting to conjecture that the TEN remains sensitive to the spins of the underlying deconfined anyon excitations. In a forthcoming work, we address more comprehensively the connection between TEN and the braided tensor category structure of imTO. It would likewise be interesting to understand novel decoherence induced negativity transitions [12, 24, 76] that may result from (competition between) the different decoherence channels discussed in this work.
@@
We conclude by commenting on practical implications of our work for imTO in open quantum systems. Currently, preparing states with chiral TO requires sequential quantum circuits [32] and, although unproven, it is widely believed that no finite-depth quantum circuit can disentangle such states from the surface of a 3D WW model; for instance, Ref. [31] proved that either there exists a commuting projector Hamiltonian for the 2D chiral semion TO (which Ref. [1] argued should not exist) or that the circuit that disentangles this TO from the surface of a 3D WW is not finite-depth. On the other hand, single shot measurement and feedback protocols for preparing ground states with Abelian TO have recently been proposed [43]. Our results thus open the door towards the dissipative preparation of chiral TOs using finite-depth LPQCs: given a parent TO that can be prepared using a single-shot measurement and feedback circuit, we have shown that appropriately engineering a locally correlated noise channel can lead to chiral imTO.
@@ Acknowledgements.
We are particulary grateful to Dominic Else, Tim Hsieh, Kohei Kawabata, Shinsei Ryu, Shengqi Sang, Nati Seiberg, Wilbur Shirley, Nikita Sopenko, Xiao-Qi Sun, and Chong Wang for stimulating discussions. R.S. is especially grateful to Shinsei Ryu for pointing out, and discussions regarding, Ref. [54]. A. P. thanks Fiona Burnell, Sanket Chirame, and Sarang Gopalakrishnan for collaboration on a related project. This material is based upon work supported by a Simons Investigator Grant from the Simons Foundation (Grant No. 566116) awarded to Shinsei Ryu (R. S.), the Sivian Fund at the Institute for Advanced Study, and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A. P.).

### 22 Mar 2024 v2

@@ I Introduction
Recent years have witnessed remarkable progress in preparing and manipulating such states in programmable quantum simulators [6, 7, 8, 9, 10]. Decoherence is invariably present in these platforms and thus identifying a sharp notion of mixed-state topological order is not merely of fundamental interest, but also of immediate practical import. While any finite temperature is known to destroy topological order (TO) in two spatial dimensions (2D) [11, 12], for local decoherence below a certain threshold, the quantum information encoded in a topological QECC is recoverable. Indeed, the persistence and eventual breakdown of topological order in a pure state |$\ket{\psi}$ subject to a local decoherence channel $\mathcal{E}$ has recently been studied through the lens of the entanglement properties of the “corrupted” density matrix [|$\mathcal{E}[\ket{\psi}\bra{\psi}]$ [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. While naïvely one expects that local errors destroy quantum correlations (and hence TO), the decohered state is not the Gibbs state and can in principle encode structured entanglement. Strikingly, for initial pure symmetry protected topological (SPT) phases, the decohered density matrix can display intrinsically mixed SPT order [23, 24, 25, 26]. While Ref. [27] recently observed a non-trivial topological entanglement negativity in a decohered Toric code (indicating imTO), a systematic framework for addressing intrinsically mixed-state topological order (imTO) is lacking.
@@
The balance of this paper is organized as follows: in Sec. II, we discuss the set of local decoherence channels under consideration and show that, for maximal decoherence, the resulting mixed-state belongs to the code space of a TSSC. In particular, we show that decoherence provides a physical mechanism for “gauging out” certain anyons, whereby only those anyons which braid trivially with the decohered anyons remain as deconfined excitations in the resulting theory. We illustrate this framework through examples in Sec. III, where we show that the decohered state can host a quantum memory as well as chiral or even non-modular topological order. In Sec. IV, we argue that our framework can naturally be generalized to include parent non-Abelian theories, which leads to our claim that braided tensor categories provide a partial classification for imTO. We then state our results in the language of strong 1-form symmetries in Sec. V, where we also show an analogy between imTOs and surface states of Walker-Wang models. Finally, we conclude in Sec. VI with a discussion of open questions and future directions.
@@ III.3 Chiral Semion from Double Semion
As an instance of this mapping, let us take the doubled semion anyon theory as our parent TO. This theory can be realized as a Pauli stabilizer Hamiltonian [32] and its anyons form a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ group under fusion, with elements $\{1,s,\bar{s},s\bar{s}\}$. Here, $s$ is a semion (it has self-statistics $\theta(s)=i$), $\bar{s}$ is an anti-semion ($\theta(\bar{s})=-i$), and $s\bar{s}$ is a boson. Now, we subject a ground state of this system to an error channel that incoherently proliferates the semion $s$ i.e., $\hat{\mathcal{A}}=\{1,s\}$. This corresponds to gauging out $s$, which braids trivially with $\bar{s}$. The resulting anyon theory for the decohered mixed-state is given by $\mathcal{A}=\{1,\bar{s}\}$ i.e., it is the chiral (anti)-semion Abelian anyon theory. This zero-correlation length state represents an intrinsically mixed-state phase of matter as chiral UMTCs are widely held to be forbidden in the ground states of locally commuting Hamiltonians [46] and are not expected to be realized in fixed point wave-functions. Since $\bar{s}$ has non-trivial self-statistics, the decohered code space encodes exactly one logical qubit; this example represents the minimal model in which one obtains a chiral anyon theory with an encoded logical qubit.
@@ IV Decoherence as gauging out in general UMTCs
IV.1 $\mathbb{Z}_{2}^{(1)}$ TSSC from Chiral Ising UMTC
@@ IV.3 Non-modular imTOs from Doubled $SU(2)_{k}$ UMTC
Thus, we can obtain a large family of imTOs by exposing the ground states of string-net models to local error channels, where the decohered code space generically retains logical information i.e., it is a decoherence-free subspace. The presence of non-trivial logical information (or a quantum memory) is encoded in the Wilson-loop algebra (equivalently, the $S$-matrix of $\mathcal{A}$). Since we have shown that the resulting imTOs can host transparent anyons, we obtain a partial classification of imTO in terms of braided fusion categories.
@@ VI Discussion
A pressing issue is to characterize our family of imTOs via their entanglement structure. While the entanglement entropy has previously been studied in mixed-state TO [51, 66], a more natural probe of entanglement in mixed-states is provided by the entanglement negativity which, unlike the entanglement entropy, is a good measure of quantum correlations in a mixed state [67, 68, 69, 70, 71]. In pure state TO, the negativity has been shown to receive universal contributions which are sensitive to the modular data of the TO (namely, the total quantum dimension) [72, 73, 74, 75, 76, 77, 12]. This topological entanglement negativity (TEN) has also been shown to be sensitive to the breakdown of TO in thermal states [78, 12], which the entanglement entropy does not accurately reflect. Since we have shown that imTO is generally characterized by non-modular anyon theories, it is an intriguing question as to what universal data the TEN captures in these states. In one specific instance, Ref. [27] distinguished between the $\mathbb{Z}_{2}^{(0)}$ and $\mathbb{Z}_{2}^{(1)}$ imTOs (obtained via decoherence of the $\mathbb{Z}_{2}$ Toric code) by the respective absence and presence of topological contributions to the negativity. Recalling that these two states correspond to quantum condensates of bosonic and fermionic loops, it is tempting to conjecture that the TEN remains sensitive to the spins of the underlying deconfined anyon excitations. In a forthcoming work, we address more comprehensively the connection between TEN and the braided tensor category structure of imTO. It would likewise be interesting to understand novel decoherence induced negativity transitions [15, 27, 79, 80] that may result from (competition between) the different decoherence channels discussed in this work.
@@
We conclude by commenting on practical implications of our work for imTO in open quantum systems. Currently, preparing states with chiral TO requires sequential quantum circuits [35] and, although unproven, it is widely believed that no finite-depth quantum circuit can disentangle such states from the surface of a 3D WW model; for instance, Ref. [34] proved that either there exists a commuting projector Hamiltonian for the 2D chiral semion TO (which Ref. [1] argued should not exist) or that the circuit that disentangles this TO from the surface of a 3D WW is not finite-depth. On the other hand, single shot measurement and feedback protocols for preparing ground states with Abelian TO have recently been proposed [47]. Our results thus open the door towards the dissipative preparation of chiral TOs using finite-depth LPQCs: given a parent TO that can be prepared using a single-shot measurement and feedback circuit, we have shown that appropriately engineering a locally correlated noise channel can lead to chiral imTO. Surprisingly, since the doubled state in our construction can always be represented as a fixed point projected entangled pair state (PEPS) with finite bond dimension as we have a topological stabilizer code in the doubled space, our work also indicates the existence of a fixed point projected entangled pair operator (PEPO) representation for density matrices exhibiting chiral topological order (with finite bond dimension). This is an intriguing implication, as it is widely believed—though not proven—that there do not exist exact PEPS representations for (interacting) chiral topological pure states.
@@ Acknowledgements.
We are particulary grateful to Dominic Else, Tarun Grover, Tim Hsieh, Kohei Kawabata, Shinsei Ryu, Shengqi Sang, Nati Seiberg, Wilbur Shirley, Nikita Sopenko, Xiao-Qi Sun, Ruben Verresen, and Chong Wang for stimulating discussions. R.S. is especially grateful to Shinsei Ryu for pointing out, and discussions regarding, Ref. [58]. A. P. thanks Fiona Burnell, Sanket Chirame, and Sarang Gopalakrishnan for collaboration on a related project. This material is based upon work supported by a Simons Investigator Grant from the Simons Foundation (Grant No. 566116) awarded to Shinsei Ryu (R. S.), the Sivian Fund at the Institute for Advanced Study, and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A. P.).